# Distributionally Robust Acceleration Control Barrier Filter for Efficient UAV Obstacle Avoidance

### v1

@@ III-B Obstacle Risk Allocation
Risk allocation helps to select the appropriate response in case of multiple obstacles. We allocate a global risk budget $\alpha_{\mathrm{total}}\in(0,1)$ among $N$ obstacles indexed by $i,j\in\{1,2,\ldots,N\}$. We assign obstacle budgets $\alpha_{i}$ using a proximity-velocity weighting with $v_{i}^{\mathrm{cl}}$ (4) for some small $\varepsilon>0$ to ensure numerical safety:
@@
The $10^{-3}$ term in $w_{i}$ prevents division by zero. The per-obstacle budgets sum to the global risk, $\sum_{i}\alpha_{i}=\alpha_{\mathrm{total}}$. The $r_{i}$ is a normalized closing-speed score that prioritizes obstacles approaching faster than others, and $w_{i}$ is the combined weight for approach urgency and proximity. Therefore, the provided obstacle budgets $\alpha_{i}$ with a one-sided Cantelli inequality [Cantelli] yields the tightening factor $\lambda_{i}$
@@
which allows to tighten the half-space right-hand. Substituting into (16) gives the tightened constraint:
@@ III-C Obstacle Acceleration Estimation
In addition to risk allocation, DOA needs to estimate acceleration bounds for the obstacle to maintain a safe distance in all cases. In order to assess the unknown acceleration of an obstacle $\hat{a}_{\max,i}$, we assume piecewise constant relative acceleration over the lookahead $H$ and compute the upper bound $\hat{a}_{\max,i}$ via higher-order SMD [2nSDM, SDM]. From measured obstacle velocity $v_{i}$, the SMD internal states $(z_{0},z_{1},z_{2})$ yield a robust estimate $\hat{a}_{i}=z_{1}$ and (if required) an estimate $\hat{j}_{i}=z_{2}$. The differentiator dynamics are given as
$\begin{split}\dot{z}_{0}&=-l_{1}\,|z_{0}-v_{i}|^{1/2}\,\mathrm{sgn}(z_{0}-v_{i})+z_{1}\\ \dot{z}_{1}&=-l_{2}\,\mathrm{sgn}(z_{0}-v_{i})\\ \dot{z}_{2}&=-\,l_{3}\,\mathrm{sgn}\big(z_{2}-\dot{z}_{1}\big)\end{split}$ (24)
where $(z_{0},z_{1},z_{2})$ are internal states and $l_{1},l_{2},l_{3}>0$ are design gains. The SMD output provides a robust, noise-insensitive estimate of the true derivative $\dot{v}_{i}=a_{i}$ with finite-time convergence. The estimated acceleration bound is updated online as
@@ III-D DR-CVaR Early Warning
Additionally, the framework uses an early trigger to complement ACBF based on conditional value at risk (CVaR) [cvar2023, drcvar2023]. We introduce a real-time distributionally robust CVaR (DR-CVaR) layer that operates in conjunction with the proposed ACBF controller. The DR-CVaR predicts the obstacle’s future state to check if it violates a dynamic distance to the UAV. This layer does not enforce the barrier constraints directly; it only anticipates imminent violations and triggers avoidance early. Once triggered, the system hands control to the ACBF block, which operates on horizon $H$.
For each obstacle $i$ and prediction time $t_{k}\in[0,\tau]$, the set of propagated sample states is defined by $P^{(n)}(t_{k})$ and $V^{(n)}(t_{k})$ for the $n$-th sample prediction. The $\tau$ denotes the reaction latency and handles preemptive risk checks, while $H$ ensures constraint satisfaction. For each sample $P^{(n)},V^{(n)}$, we define the LoS unit vector, i.e., $n^{(n)}$ and the closing speed $v^{(n)}$ computed as in Eq. (4). This helps to estimate a dynamic distance for the $n$-th sample
@@
where $a_{\mathrm{i}}$ and $a_{\mathrm{e}}$ denote the available deceleration magnitudes of the obstacle and the UAV deceleration. Then, the safe distance is
@@
To guard against distributional mismatch, we add a Wasserstein upshift $\epsilon L_{Z}$ with radius $\epsilon$ and Lipschitz constant $L_{Z}$ [wasserstein]. We adopt the standard empirical CVaR definition using the worst $m$ samples [cvar2023, drcvar2023]
@@
where $m=\lceil\beta N_{s}\rceil$ with $\beta\in(0,1)$ specifies how much of the worst-case distribution we consider. $N_{s}$ denotes the number of Monte-Carlo samples per obstacle used for short-horizon prediction. We set $\epsilon$ as a robust radius tied to the sample spread. If $\max_{k}\mathrm{CVaR}_{\beta}^{\mathrm{DR}}(Z)\geq 0$, the early warning layer triggers avoidance before any barrier violation is imminent. It replaces the nominal acceleration $a_{\mathrm{n}}$ by an evasive avoid acceleration $a_{\mathrm{e}}$ with a projection from ACBF. It tightens the ACBF constraints and $R_{eff}$ by reducing $\alpha_{\mathrm{total}}$ and increasing $H$ for one cycle. The $\beta$ controls the tail fraction for prediction-based early warning, and the $\alpha_{\mathrm{total}}$ provides the probabilistic tightening of the ACBF constraints.
@@ III-E Gauss-Southwell Projection
To further improve the real-time capabilities, we avoid QP and apply a deterministic Gauss-Southwell (most violated constraint) projection to a nominal acceleration command $a_{0}=a_{\mathrm{n}}$ [fastproj, gauss]
@@
where $i^{\star}$ is the index of the most violated constraint and $a_{l}$ is the acceleration iterate at the current step of the projection loop. The projection is iterated for a fixed number of steps. At each iteration $l$, the acceleration is updated as
@@
where $\omega\in(0,1]$ is a relaxation coefficient in the projection algorithm, which is tuneable empirically for fastest convergence. The projection loop runs for a fixed iteration budget, which is typically limited to 15 iterations. The resulting $a_{e}=a_{l}$ is the safe acceleration command. The iteration (31) is a relaxed Gauss–Southwell projection onto the most violated halfspace, followed by projection onto the acceleration space (16). This scheme is known to be numerically stable [fastproj, gauss] and iterates toward the feasible set $C$ (8).
@@ IV Simulation Study
In this section, we compare our DR-ACBF with three other collision avoidance methods in a simulation study. We use the classical CBF-QP [ames2014acc, AdaptiveCBF2022], robust policy CBF (RP-CBF) [SafetyFly2025], and model predictive control (MPC) [lindqvist2020ral, ampc, tinyMPC] as baselines and integrate the DR-CVaR early trigger for a fair comparison of the ACBF improvement.
We use the following metrics for our comparison: The success rate $s$ is defined as the ratio of simulation runs that successfully avoid collisions over all runs. The computation time $t_{cp}$ is the processing time of our avoidance algorithm (from acceleration estimation to projection in Figure 2). The command time $t_{cm}$ represents the control latency, defined as the time from the trigger to issuing the first avoid command. The reaction time $t_{r}$ represents the duration from the trigger to the first detectable avoid action and combines the command time with the dynamics of the UAV. The minimum separation $d_{s}$ represents the minimum distance between the UAV and the obstacle during the avoidance maneuver. Finally, $v_{cl}$ represents the speed of the obstacle towards the UAV.
@@ IV-A Simulation Setup
DR-ACBF uses a one-step global collision risk budget of $\alpha_{\mathrm{total}}=0.10$, distributed among obstacles according to proximity $d_{i}$ and closing speed $v^{\mathrm{cl}}_{i}$. A single horizon per cycle $H$ is selected via coarse search $H\in[H_{\min},H_{\mathrm{max}}]$ with user given time horizon $H_{\mathrm{max}}=0.4$ s and $H_{\min}=0.15$. The coarse search selects the largest feasible $H$, which maximizes clearance while still satisfying all constraints. Obstacle acceleration envelopes $\hat{a}_{\max,i}$ are learned online using an SMD differentiator with $(\gamma,L_{0},L_{1},L_{2})=(1.5,4,3,2)$. Safe accelerations are obtained via a constant-time Gauss–Southwell projector with 12 base iterations and relaxation $\omega=1.0$.
The complete DOA consists of the avoidance action, followed by a hover at the bypass position, and mission continuation towards the next target position. The UAV should hover with time to stop within $t_{\mathrm{s}}=0.5$ s and braking intensity of $k_{\mathrm{s}}=1.0$. The lateral clearance $d_{cl}$ is $0.05$ m, and minimum hold time after avoid $t_{hd}$ is 3 s. The DR-CVaR early-warning trigger runs before projection with $\beta=0.05$ and a reaction latency of $\tau=0.02$ s. In our experiments, we set the Wasserstein radius $\epsilon$ to 0.05 for numerical stability.
@@ IV-B Simulation Results
Table I compares the proposed DR-ACBF approach against conventional CBF, MPC, and recent RP-CBF techniques with identical simulation scenarios of one or two dynamic obstacles. Our DR-ACBF outperforms the other techniques in all time measurements. Particularly, the processing time is reduced by 19% to 42%, which is important for embedded platforms. Despite the reduced computational effort, DR-ACBF maintains comparable performance with a mean success rate $s$ of 84.85% and the minimum separation $d_{s}$ of 1.844 m, satisfying safety constraints throughout the mission.
@@
MPC 1 86.0 7.13 16.00 6.00 - 0.909
2 83.4 9.80 13.56 3.08 - 3.040
Mean 84.7 8.36 14.85 4.65 0.591 1.888
CBF-QP 1 86.1 7.14 25.25 3.15 - 0.913
2 83.6 9.82 12.95 2.31 - 3.098
Mean 84.9 8.36 19.62 2.77 0.461 1.914
RP-CBF 1 86.0 7.13 15.35 2.68 - 0.908
2 83.4 9.80 23.94 2.54 - 3.040
Mean 84.77 8.36 19.29 2.62 0.425 1.844
DR-ACBF 1 86.1 7.14 12.29 1.68 - 0.913
2 83.5 9.65 14.08 1.58 - 2.990
Mean 84.85 8.30 11.98 1.64 0.347 1.844
@@ IV-C Sensitivity Analysis and Ablation
To test the robustness of the proposed method, we conducted a sensitivity analysis using 1000 simulation runs and varying three key parameters. Table II summarizes the sensitivity of DR-ACBF to the key parameters maximum time horizon $H_{\mathrm{max}}$, clearance distance $d_{cl}$, and hold time $t_{hd}$. Each parameter was varied independently while keeping the others fixed. The key parameters of the first row in Table II correspond to the simulation results over 3000 runs (Table I). Results indicate that DR-ACBF maintains a consistent $s$ above 77% and shows no significant change in $t_{r}$ and $d_{s}$ across all key parameter configurations. Minor variations in $H_{\mathrm{max}}$ shows a reduction in $t_{cm}$ and $t_{cp}$.
@@
OFF 1 32.00 9.70 11.74 1.74 0.707
2 37.90 8.01 12.06 2.06 0.858
Mean 34.90 8.70 11.93 1.93 0.796
Finally, we conduct an ablation study to evaluate the effectiveness of the proposed early trigger DR-CVaR in ACBF. In order to isolate the contribution of the proposed DR-CVaR trigger, we compared it with a baseline, i.e., a minimal trigger that uses a single-time Gaussian chance constraint at the time of closest approach (DR-CVaR OFF in Table III). All other components (geometry, dynamics, parameters) remain identical across 1000 runs. As shown in Table III, activating DR-CVaR leads to a substantial improvement in performance, increasing the success rate $s$ from 34.9% to 82.5% while maintaining comparable response times. The separation distance $d_{s}$ improves from 0.796 m to 1.881 m, indicating enhanced safety margins and more reliable avoidance behavior. These results clearly demonstrate that the DR-CVaR trigger is essential for improving safety. Note that the success rate of ACBF can also be improved with DR-CVaR OFF by increasing both $H_{\min}$ and $H_{\max}$, but at the cost of slower reaction time. Thus, the DR-CVaR trigger enables fast reaction.

### v2

@@ III-B Obstacle Risk Allocation
Risk allocation helps to select the appropriate response when multiple obstacles are present by avoiding uniform conservatism and prioritizing the nearest, fastest-closing obstacle. We allocate a global risk budget $\alpha_{\mathrm{total}}\in(0,1)$ among $N$ obstacles indexed by $i,j\in\{1,2,\ldots,N\}$. We assign obstacle budgets $\alpha_{i}$ using a proximity-velocity weighting with $v_{i}^{\mathrm{cl}}$ (4) for some small $\varepsilon>0$ to regularize the normalization and ensure numerical safety:
@@
The $10^{-3}$ term in $w_{i}$ prevents division by zero. The per-obstacle budgets sum to the global risk $\sum_{i}\alpha_{i}=\alpha_{\mathrm{total}}$. The $r_{i}$ is a normalized closing-speed score that prioritizes obstacles approaching faster than others, and $w_{i}$ is the combined weight for approach urgency and proximity. Therefore, the provided obstacle budgets $\alpha_{i}$ with a one-sided Cantelli inequality [27] yields the tightening factor $\lambda_{i}$
@@
which allows to tighten the half-space right-hand. The quantity $d_{\text{true},i}(t+H)$ denotes the unknown realized distance at time $t+H$ which is modeled around $d_{\text{pred},i}(t+H)$ for sensing uncertainty as a random variable with standard deviation $\sigma_{dH,i}$ to predict the LoS distance. Substituting the uncertainty safety margin into (15) gives the tightened constraint:
@@ III-C Obstacle Acceleration Estimation
In addition to risk allocation, DOA needs to estimate acceleration bounds for the obstacle to maintain a safe distance in all cases. In order to assess the unknown acceleration of an obstacle $\hat{a}_{\max,i}$, we assume piecewise constant relative acceleration over the lookahead $H$ and compute the upper bound $\hat{a}_{\max,i}$ via higher-order SMD [8, 33]. From measured obstacle velocity $v_{i}$, the SMD internal states $(z_{0},z_{1})$ yield a robust estimate $\hat{a}_{i}=z_{1}$. The differentiator dynamics are given as
[…]
where $(z_{0},z_{1})$ are internal states and $l_{1},l_{2}>0$ are design gains. The SMD output provides a robust, noise-insensitive estimate of the true derivative $\dot{v}_{i}=a_{i}$ with finite-time convergence. The estimated acceleration bound is updated online as
@@ III-D DR-CVaR Early Warning
Additionally, the framework uses an early trigger to complement ACBF based on conditional value at risk (CVaR) [24, 31]. We introduce a real-time distributionally robust CVaR (DR-CVaR) layer that operates in conjunction with the proposed ACBF controller. The DR-CVaR trigger samples predictions from the tracker’s Gaussian output, assuming constant acceleration and bounded braking actions to check if it violates a dynamic distance to the UAV. The DR-CVaR anticipates imminent violations and triggers early avoidance. It acts as an early-warning layer that, when activated, tightens ACBF enforcement by adjusting horizon and risk, and only triggers when the risk condition is met (see Fig. 2). Once triggered, the system hands control to the ACBF block, which operates on horizon $H$. For each obstacle $i$ and prediction time $t_{k}$ sampled on a discrete grid within $0,H_{max}]$, the set of propagated sample states is defined by $P^{(n)}(t_{k})$ and $V^{(n)}(t_{k})$ for the $n$-th sample prediction. For each sample $P^{(n)},V^{(n)}$, we define the LoS unit vector, i.e., $n^{(n)}$ and the closing speed $v^{(n)}$ computed as in Eq. (4). This helps to estimate a dynamic distance for the $n$-th sample
@@
where $a_{\mathrm{i}}$ and $a_{\mathrm{e}}$ denote the available magnitudes of the obstacle and the UAV deceleration. The $\tau$ denotes the reaction latency. Then, the safe distance is
@@
To guard against distributional mismatch, we add a Wasserstein upshift $\epsilon L_{Z}$ with radius $\epsilon$ and Lipschitz constant $L_{Z}$ [5]. We adopt the standard empirical CVaR definition using the worst $m$ samples [24, 31]
@@
where $m=\lceil\beta N_{s}\rceil$ with $\beta\in(0,1)$ specifies how much of the worst-case distribution we consider. $N_{s}$ denotes the number of Monte-Carlo samples per obstacle used for short-horizon prediction. We set $\epsilon$ as a robust radius tied to the sample spread. If $\max_{k}\mathrm{CVaR}_{\beta}^{\mathrm{DR}}(Z)\geq 0$, the early warning layer triggers avoidance before any barrier violation is imminent. The trigger activates ACBF with a projection to replace the nominal acceleration $a_{\mathrm{n}}$ by an evasive avoid acceleration $a_{\mathrm{e}}$. It tightens the ACBF constraints and $R_{eff}$ by reducing $\alpha_{\mathrm{total}}$ and increasing $H$ for one cycle. The $\beta$ controls the tail fraction for prediction-based early warning, and the $\alpha_{\mathrm{total}}$ provides the probabilistic tightening of the ACBF constraints.
@@ III-E Gauss-Southwell Projection
To improve the real-time capabilities, we avoid QP and apply a deterministic Gauss-Southwell (most violated constraint) projection to a nominal acceleration $a_{0}=a_{\mathrm{n}}$ [12, 26]
@@
where $i^{\star}$ is the index of the most violated constraint and $a_{l}$ is the acceleration iterate at the current step of the projection loop. At each iteration $l$, the acceleration is updated as
@@
where $\omega\in(0,1]$ is a relaxation coefficient in the projection algorithm, which is tuneable empirically for fastest convergence. The resulting $a_{e}=a_{l}$ is the safe acceleration command. The iteration (30) is a relaxed Gauss-Southwell projection onto the most violated halfspace, followed by projection onto the acceleration space (15). In addition, the loop can stop early when no constraint is violated, i.e., when $\max_{i}(A_{i}a_{l}-b_{i})\leq 0$. Gauss-Southwell projection is known to be numerically stable [12, 26] and fast by limiting the iterations. The projector iterates to satisfy (22) and reach the feasible set $C$ (8). We run a fixed number of iterations and check the maximum violation; if it exceeds a threshold, a fallback is triggered by increasing $H$ and tightening the risk budget. If violations persist, the projector biases the command toward braking and landing.
@@ IV Simulation Study
In this section, we use the classical velocity-based CBF-QP [2, 34], robust policy CBF (RP-CBF) [14], and model predictive control (MPC) [16, 11, 25] as baselines and integrate the identical DR-CVaR early trigger on all approaches for a fair comparison of the ACBF improvement. For the MPC baseline [20, 11], we use MATLAB’s built-in MPC (QP solver) with prediction horizon $p=20$ and control horizon $m=3$ for real-time constraints. For RP-CBF [14], horizon $T_{\text{pcbf}}=3.0$ s and $N=32$ disturbance samples are incorporated via tracker’s obstacle-motion estimation. The CBF-QP uses the RP-CBF configuration and horizon of $0.4$ s.
We use the following metrics for our comparison: the success rate $s$ is defined as the ratio of simulation runs that successfully avoid collisions over all runs. The computation time $t_{cp}$ is the processing time of our avoidance algorithm (from acceleration estimation to projection in Figure 2). The command time $t_{cm}$ represents the control latency, defined as the time from the trigger to issuing the first avoid command. The reaction time $t_{r}$ represents the duration from the trigger to the first deviation from the nominal command and combines the $t_{cm}$ with the dynamics of the UAV. The minimum separation $d_{s}$ represents the minimum distance between the UAV and the obstacle during the avoidance maneuver. Finally, $v_{cl}$ represents the speed of the obstacle towards the UAV.
@@ IV-A Simulation Setup
DR-ACBF uses a one-step global collision risk budget of $\alpha_{\mathrm{total}}=0.10$, distributed among obstacles according to proximity $d_{i}$ and closing speed $v^{\mathrm{cl}}_{i}$. A single horizon per cycle $H$ is selected via coarse search $H\in[H_{\min},H_{\mathrm{max}}]$ with user given time horizon $H_{\mathrm{max}}=0.4$ s and $H_{\min}=0.15$ s. The coarse search selects a feasible $H$ that maximizes clearance and satisfies all constraints. Obstacle acceleration envelopes $\hat{a}_{\max,i}$ are learned online using an SMD differentiator with $(\gamma,L_{0},L_{1})=(1.5,4,3)$. Safe accelerations are obtained via a fixed-time Gauss-Southwell projector with 15 iterations and relaxation $\omega=1.0$. The complete DOA consists of the avoidance action, followed by a hover at the bypass position, and mission continuation towards the next target position. For avoidance direction, the best of 12 candidate directions (axis/diagonal) for each step is selected to minimize lateral acceleration over $H$, providing a bypass position with clearance offset. The UAV should hover with time to stop within $t_{\mathrm{s}}=0.5$ s and braking intensity of $k_{\mathrm{s}}=1.0$. The lateral clearance $d_{cl}$ is $1.0$ m, and minimum hold time after avoid $t_{hd}$ is 0.2 s. The DR-CVaR early-warning trigger runs before projection with $\beta=0.05$ and a reaction latency of $\tau=0.02$ s. We set the Wasserstein radius $\epsilon$ to 0.05 for numerical stability.
@@ IV-B Simulation Results
This section presents the results for the first scenario, where the trajectory of one obstacle is curving to intersect the UAV’s path, and the other obstacle moves directly towards the UAV. Both obstacles start at 50 m from the UAV with small detection noise and increasing speed. Table I compares the proposed DR-ACBF approach against conventional CBF, MPC, and recent RP-CBF techniques with identical simulation scenarios of two dynamic obstacles. Our DR-ACBF outperforms the other techniques in all time measurements. Particularly, the processing time is reduced by 31% to 54% with lower variability, which is important for embedded platforms. Despite the reduced computational effort, DR-ACBF maintains comparable performance with a mean success rate $s$ of 99.64% and the minimum separation $d_{s}$ of 4.73 m, satisfying safety constraints throughout the mission. Standard deviations reflect variability for methods (with extreme outliers removed); overall low for DR-ACBF and high $t_{r}$ RP-CBF’s due to obstacle 1.
@@
MPC 1 99.52 10.43 10.14 3.23 - 2.82
2 99.89 10.97 10.00 3.21 - 6.64
Mean 99.71 10.70 10.07 3.22 0.956 4.73
Std - - 4.52 0.93 0.666 -
CBF-QP 1 99.51 10.43 10.03 2.21 - 2.82
2 99.84 11.00 10.00 2.19 - 6.65
Mean 99.67 10.71 10.02 2.20 0.714 4.74
Std - - 1.31 0.46 0.476 -
RP-CBF 1 99.52 10.43 106.26 1.87 - 2.82
2 99.89 10.97 13.39 2.26 - 6.64
Mean 99.71 10.70 59.83 2.08 0.642 4.73
Std - - 93.48 0.48 0.416 -
DR-ACBF 1 99.48 10.43 10.06 1.39 - 2.82
2 99.80 10.98 10.00 1.41 - 6.64
Mean 99.64 10.70 10.03 1.40 0.444 4.73
Std - - 2.76 0.38 0.270 -
@@ IV-C Sensitivity Analysis and Ablation
We assessed the robustness of the proposed method and the tracker’s impact via a sensitivity analysis and ablation study with 1000 runs in a second scenario, varying three key parameters. The second scenario has one obstacle crossing diagonally and another moving straight toward the UAV, with initial distances of 150 m and 200 m. This scenario tests linear-motion avoidance under higher long-range detection noise, affecting tracking and reducing success. Table II summarizes the sensitivity of DR-ACBF to the key parameters maximum time horizon $H_{\mathrm{max}}$, clearance distance $d_{cl}$, and hold time $t_{hd}$. Each parameter was varied independently while keeping the others fixed. Results indicate that DR-ACBF maintains a consistent $s$ above 77% in this scene and shows no significant change in $t_{r}$ and $d_{s}$ across all key parameter configurations. Minor variations in $H_{\mathrm{max}}$ show a reduction in $t_{cm}$ and $t_{cp}$.
@@
Finally, we conduct an ablation study to evaluate the effectiveness of the proposed early trigger DR-CVaR in ACBF. We compared it with a baseline, i.e., a simple trigger that uses a single-time Gaussian chance constraint by evaluating relative distance at that single time (DR-CVaR OFF in Table III). The simple trigger is required as the ACBF behaves like a reactive safety filter; without the trigger, it may activate late under latency, even with deterministic settings. All other components (geometry, dynamics and parameters as in first row of Table II) remain identical across 1000 runs. Table III shows that DR-CVaR ON leads to a substantial improvement in performance, increasing the success rate $s$ from 34.9% to 82.5% while maintaining comparable response times. This highlights the need for an early trigger; the adaptive DR-CVaR ensures timely ACBF activation, whereas a simple trigger leads to late or no ACBF activation, significantly reducing the success rate.
@@
OFF 1 32.0 9.70 11.74 1.74 0.707
2 37.9 8.01 12.06 2.06 0.858
Mean 34.9 8.70 11.93 1.93 0.796
The separation distance $d_{s}$ improves from 0.796 m to 1.881 m, indicating enhanced safety margins and more reliable avoidance behavior. These results clearly demonstrate that the DR-CVaR trigger is essential for improving safety. Note that the success rate of ACBF can also be improved with DR-CVaR OFF by increasing both $H_{\min}$ and $H_{\max}$, but at the cost of slower reaction time. Thus, the DR-CVaR trigger enables fast reaction. We compared projection-based and QP-based DR-ACBF with the same setup. Success rates were similar, but QP-based had larger mean $t_{cp}$ and $t_{cm}$ (0.766 ms, 5.36 ms) than projection-based DR-ACBF (0.475 ms, 2.24 ms).